OpenStudy (anonymous):

2(–2 + 2x) = 4(x – 3)

OpenStudy (imtiaz7):

Its simple multiplication: -4+4x=4x-12 -4+4x-4x+12=0 (Here +4x and -4x will be cancelled) -4+12=0 8